OpenStudy (anonymous):

Sally travels by car from one city to another. She drives for 30.0 min at 80.0 km/h, 12.0 min at 105 km/hr, and 45.0 min at 40.0 km/h, and she spends 15.0 min eating lunch and buying gas. Determine the average speed and the total distance for the trip.

OpenStudy (anonymous):

first we change all the times from min to hours dividing by 60 \[30 \min = 0.5 h; 12 \min = 0.2 h; 45\min =0.75h; 15\min =0.25h \]she traveled \[time \times speed =\Delta x\]\[0.5h \times 80km/h = 40km\]\[0.2h \times 105km/h =21km\]\[0.75h \times40km/h =30 km\]\[0.25h \times0km/h =0km\]total distance for the trip was 40+21+30+0=91km and the average speed is:\[\frac{ \Delta x }{ \Delta t }=\frac{ 91km }{ (0.5+0.2+0.75+0.25) h }=\frac{ 91km }{ 1.7h }=5.352km/h\]

OpenStudy (anonymous):

Thank you!